Assume two consumers A and B have the same utility function U= x1.x2. Prices are p1=2 , p2=5 ,income = 100 (for each) initial endowments are given as Wa= (30,8) and Wb=(25, 10) . Derive excess demand
1
Expert's answer
2017-04-17T09:38:05-0400
Consumers A and B have the same utility function U = x1*x2, p1 = 2 , p2 = 5, I = 100, initial endowments are Wa = (30, 8) and Wb = (25, 10).
Excess demand for product 1 is 30 + 25 = 55 and for product 2 is 8 + 10 = 18.